# Stackelberg vs. Nash in the Lottery Colonel Blotto Game

### v1

@@ Appendix C OMITTED PROOFS FROM SECTION 5
Previous works Kim et al. [2018]; Li and Zheng [2022] have established that if the strategy profile $(\bm{x}_{a},\bm{x}_{b})$ is a NE, then $x_{ij}>0$, for $\forall i\in\{a,b\}$, $\forall j\in[n]$. Hence, if $(\bm{x}_{a},BR(\bm{x}_{a}))$ constitutes a NE, we must have $K(\bm{x}_{a})=[n]$. Therefore, we examine the situation where the leader’s commitment results in the follower having positive best responses across all battlefields.
@@
$\frac{x_{a}}{x_{b}}=\frac{t_{4}}{t_{3}}.$ (34)
@@
It is worth noting that when $\frac{v_{aj}}{v_{bj}}=c$ where $c$ is a constant, the case (1) is equivalent to case (2), and $t_{4}=t_{3}=0$. ∎

### v3

@@ Appendix C OMITTED PROOFS FROM SECTION 5
Previous works have established that if the strategy profile $(\bm{x}_{a},\bm{x}_{b})$ is a NE, then $x_{ij}>0$, for $\forall i\in\{a,b\}$, $\forall j\in[n]$ [Kim et al., 2018; Li and Zheng, 2022]. Hence, if $(\bm{x}_{a},BR(\bm{x}_{a}))$ constitutes a NE, we must have $K(\bm{x}_{a})=[n]$. Therefore, we examine the situation where the leader’s commitment results in the follower having positive best responses across all battlefields.
@@
First, for simplicity, let’s define a few expressions.
Definition 2.
For any $T\subseteq[n]$, $i\in\{a,b\}$, let
$v_{iT}$ $=\sum_{j\in T}v_{ij},$
$t_{1}(T)$ $=\Psi_{1}(v_{aT},v_{a\overline{T}},v_{bT},v_{b\overline{T}}),$
$t_{2}(T)$ $=\Psi_{1}(v_{a\overline{T}},v_{aT},v_{b\overline{T}},v_{bT}),$
$t_{3}(T)$ $=\Psi_{2}(v_{a\overline{T}},v_{aT},v_{b\overline{T}},v_{bT},t_{2}(T),t_{1}(T))+\Psi_{3}(v_{aT},v_{a\overline{T}},v_{bT},v_{b\overline{T}},t_{1}(T),t_{2}(T)),$
$f(v_{aT},v_{bT},v_{a\overline{T}},v_{b\overline{T}})$ $=\frac{t_{4}(T)}{t_{3}(T)}$
@@
$\frac{x_{a}}{x_{b}}=\frac{t_{4}(M)}{t_{3}(M)}=f(v_{aT},v_{bT},v_{a\overline{T}},v_{b\overline{T}})$ (36)
@@
It is worth noting that when $\frac{v_{aj}}{v_{bj}}=c$, a constant, case (1) is equivalent to case (2), and $t_{4}(M)=t_{3}(M)=0$. ∎